OpenStudy (anonymous):

The volumes of two similar cylinders are 2197pi in cubed and 8000pi in cubed find the following:

OpenStudy (anonymous):

\[\frac{a^3}{b^3} = \frac{2197 {\pi}^3}{8000{\pi}^3}\] Tell me this first: \[\large \sqrt[3]{2197{\pi}^3} = ??\] Since there is \({\pi}^3\) here so you can directly take it out of cube root brackets as \(\pi\).

OpenStudy (anonymous):

 \[\frac{ Volume I }{ Volume II }=\frac{a ^{3} }{ b ^{3} }\] \[\frac{ r1 }{ r2 }=\frac{ h1 }{ h2 }=\frac{ a }{ b }\] \[\frac{ Area I }{ Area II }=\frac{ a ^{2} }{ b ^{2} }\]